OpenStudy (anonymous):

Show that 0.02143 (line over 143) is a rational number by writing it as a fraction of integers.

OpenStudy (anonymous):

well if .5=1/2=5/10 .25=1/4=25/100 i think the pattern is 0.02143 2143/100000